OpenStudy (anonymous):

Sam is flying a kite. The length of the kite string is 80 meters, and it makes an angle of 75° with the ground. The height of the kite from the ground is (A) 20.27 (B) 61 (C)77.27 meters HELPP for a metal and fan

OpenStudy (anonymous):

@ZairaHernandez

OpenStudy (anonymous):

ANyone lol

OpenStudy (anonymous):

work hard maybe youll get it right(: lol

OpenStudy (anonymous):

I mean i ammm.. but idkk it.. helpp

OpenStudy (anonymous):

well school over im out...

OpenStudy (anonymous):

maybe tommrow

OpenStudy (anonymous):

yes bye

OpenStudy (anonymous):

 \[\sin \theta = \frac{ Opposite }{ Hypotenuse }\]

OpenStudy (anonymous):

Hold on .....

OpenStudy (anonymous):

So... \[Hypotenuse \sin \theta = Opposite\]

OpenStudy (anonymous):

Plugging in the numbers... \[80Sin(75)=77.27 Meters\]

OpenStudy (anonymous):

There's the answer

OpenStudy (anonymous):

thanksss.. I reward yuh with a medal